# GraCoRe: Benchmarking Graph Comprehension and Complex Reasoning in Large Language Models

### v1

@@ 3.3 Data Statistics
Table 2 shows several key statistics of our GraCoRe benchmark. We categorized the dataset based on graph structure into two main datasets: PureGra and HeterGra. Each main dataset contains multiple sub-datasets used for corresponding task testing. In total, there are 19 tasks with up to 5,140 graphs. Detailed statistics for each task can be found in the Appendix A.
@@ 3.4 Evaluation
This thesis evaluates the output of large language models (LLMs) using exact match accuracy. We focus on various output types, including boolean values (e.g., graph recognition tasks), integers (e.g., path lengths), floating-point numbers (e.g., similarity and average degree), and lists of nodes (e.g., paths). For tasks with multiple possible solutions (e.g., BFS), we verify whether the output constitutes a valid answer.
Since the metrics of different GraCoRe tasks are incomparable and differently sensitive, less experienced audiences cannot easily compare and interpret results, which is also prevalent in recent LLM benchmarks like KolaYu et al. (2023). Therefore, we utilized standardized global scoresDyck et al. (2005) to evaluate the performance of LLMs on each task.
@@
where $\mu$(·) and $\sigma$(·) denote the mean and standard deviation.Next, we use the Min-Max scalingPatro and Sahu (2015) method to adjust the scores to the range of 0-100, making it easier to observe and compare the results. The final scores are presented as:
@@ 4.1 Experimental Setup
Based on the GraCoRe benchmark, we aim to investigate whether language models can understand and reason about graph structures through textual descriptions of graph data. Additionally, we explore whether designed prompts can enhance the models’ performance on graph-related tasks.
Models and Settings We evaluated a total of eight popular models on the GraCoRe benchmark, including three closed-source models and five open-source models. The closed-source models are: GPT-4o, GPT-4, and GPT-3.5Achiam et al. (2023). The open-source models are: LLama3-ins-8b, LLama2-7b-chatTouvron et al. (2023), Chatglm3-6b, Chatglm2-32k-7bDu et al. (2021), Vicuna-v1.5-7b, Vicuna-v1.5-16k-7bChiang et al. (2023) and Qwen2-7b-insBai et al. (2023). More details about these models can be found in the appendix.
Figure 4: Effect of Graph Size
@@ 4.2 Main Results
Task Dimensional Analysis Table 3 and Table 4 shows the performance of various large language models across 19 different tasks in our GraCoRe benchmark. Among all tasks, graph understanding tasks are generally less challenging for the models, while graph reasoning tasks are more difficult. This conclusion is reflected in the performance of each model. Additionally, closed-source models, such as those developed by OpenAI, consistently demonstrated superior performance. The GPT-4o model achieved the highest total score of 1419, ranking first. Among open-source models, Llama3-8b and Qwen2-7b-ins also performed exceptionally well, scoring 1054 and 1053 respectively, ranking fourth and fifth, just behind the OpenAI models. In contrast, Chatglm2-7b model performance was less satisfactory compared to other open-source models. Regarding task types, performance on reasoning tasks was not ideal, but this outcome was expected.
Additionally, the average $z$-scores indicate a significant gap between open-source models and commercial closed-source models. Only the Llama3-8b and Qwen2-7b-ins models have $z$-scores greater than 0, indicating performance above the average level. This suggests that the open-source community needs more collaboration to support and improve the latest open-source large language models. It is noteworthy that these two open-source models have graph understanding and reasoning capabilities approaching those of the GPT-3.5 model.
Ability Dimensional Analysis We further analyze Table 3 and Table 4 to evaluate the overall performance of the models from a capability perspective. We use radar charts to present the second and third layers of the three-tier taxonomy in a more intuitive manner. Left part of Figure 3 shows the performance of LLMs across four different capability dimensions at the second layer, where each dimension’s score is the average of its respective tasks. It can be observed that most models perform well in graph understanding and semantic reasoning but need improvement in structural reasoning tasks. For most open-source models, performance in heterogeneous graph understanding tasks is not very satisfactory.
Right part of Figure 3 provides a more granular view of LLMs’ performance across ten capability dimensions at the third layer. It is evident that the reasoning abilities of large language models in graph theory problems are not very strong, particularly in complex graph theory tasks. Furthermore, GPT-4 and GPT-4o exhibit robust and balanced graph processing capabilities, especially excelling in complex tasks. In contrast, other models show less balanced performance in graph processing. For instance, the Vicuna-v1.5-16k-7b model excels in understanding heterogeneous graph structures but performs poorly in other areas, suggesting that the rich semantic information may enhance the model’s graph processing capabilities.
@@ 4.3 Further Analysis
Effect of Graph Size Figure 4 illustrates the performance of the model across four tasks as the number of nodes increases. Specifically, we used two graph understanding tasks and two graph reasoning tasks to evaluate the model’s performance. The results indicate that as the number of nodes increases, the performance on each task decreases. In the graph understanding tasks, the impact is particularly significant on the average degree task. We hypothesize that the increase in nodes leads to a substantial rise in computational demands, resulting in decreased accuracy. In the graph reasoning tasks, we observed that for the shortest path problem, the performance of GPT-3.5 diminishes with an increasing number of nodes. However, for the other three open-source models, the node size appears to have minimal impact. This may be because the added nodes do not affect the search for the original path, though further experiments are needed to confirm this hypothesis.
Effect of Random Sort Since our data consists of randomly generated graphs with nodes named by numbers, the performance of node-related tasks may be affected by changes in node order. In this study, we examine the impact of text order on the understanding of graph structures by large language models by comparing random sorting with sequential sorting. The results in Table 5 indicate that the model’s performance under sequential sorting is often superior to that under random sorting, particularly in graph path reasoning tasks, where the impact is significant. This suggests that renaming nodes and ordering them sequentially can enhance model performance in path reasoning problems. However, it also highlights the model’s lack of training on graph data with random sorting.
Model NN ST BR SP HP GPT-3.5 sort 0.993 0.954 0.504 0.117 0.243 random 0.968 0.865 0.346 0.070 0.048 Qwen2-7b sort 0.876 0.874 0.396 0.165 0.115 random 0.773 0.885 0.374 0.159 0.000 Llama3-ins-6b sort 0.889 0.735 0.200 0.189 0.285 random 0.786 0.809 0.200 0.198 0.059
@@
Figure 5: Effect of Text Enhancement
Heterogeneous graphs often contain rich semantic information, which typically aids models in understanding and reasoning about graph text information. If enhanced text information, such as paper titles and abstracts in the ACMText data is removed, can large language models perform reasoning based solely on the remaining structural information? Figure 5 compares the performance of different large language models on the node classification problem. The results indicate that large language models can reason about graphs without enhanced text information, particularly the GPT-4 and GPT-4o models, which maintain strong node prediction capabilities even without text information. However, for the GPT-3.5 model, the absence of text information significantly impacts performance. For other open-source models, the presence or absence of text information appears to have minimal impact.
@@ 5 Conclusion
This paper introduces GraCoRe, a benchmark designed to evaluate the capability of large language models (LLMs) in understanding and reasoning with graph-structured data. We develop a detailed, multi-level classification system to assess model performance on graph-based tasks. Utilizing GraCoRe, we evaluate 10 prominent LLMs, revealing significant deficiencies in their graph reasoning abilities. Extensive experimental results demonstrate the effectiveness of our benchmark in measuring the performance of LLMs on graph tasks.
6 Limitations
@@
Appendix A Details on the Data Generation
@@
Our dataset mainly includes two types of graph data, heterogeneous graph data and pure graph data. In addition, based on these two types of data, we set up 19 tasks to evaluate the large model, which includes a total of 11 datasets. This diversity ensures the comprehensiveness and stability of our data.
@@
Figures 6 to 26 illustrate examples of prompts and graph data descriptions for each task. To ensure that the generated results align with our capabilities and task requirements, we concatenate the prompts for each task with the initial prompts

### v2

@@ 3.3 Data Statistics
Table 2 shows several key statistics of our GraCoRe benchmark. We categorized the dataset based on graph structure into two main datasets: PureGra and HeterGra. Each main dataset contains multiple sub-datasets used for corresponding task testing. In total, there are 19 tasks with up to 5,140 graphs. Detailed statistics for each task can be found in the Appendix C.
@@ 3.4 Evaluation
This thesis evaluates LLM outputs using exact match accuracy across various output types, including boolean values (e.g., for graph recognition tasks), integers (e.g., path lengths), floating-point numbers (e.g., similarity and average degree), and lists of nodes (e.g., paths). For tasks with multiple valid solutions (e.g., BFS), we check whether the output is a valid solution.
Standardized global scores(Dyck et al., 2005) have become the mainstream choice in fields such as educational assessment and intelligence testing. Compared to using raw scores, employing standardized scores in the multi-task evaluation of LLMs offers the following advantages:
  • Standardized scores facilitate fair comparisons across different tasks and datasets. Given the varying difficulties and sensitivities of metrics across tasks, absolute performance metrics cannot be directly compared. Standardization normalizes these scores, enabling a more meaningful assessment of model performance across diverse tasks.
  • Standardized scores reduce biases from specific tasks when ranking models. Models often display varying performance across tasks, making it difficult to gauge their strengths in graph understanding and reasoning. By normalizing scores, standardized metrics provide a more accurate and equitable basis for evaluating model capabilities across a broad range of tasks.
Since the metrics of different GraCoRe tasks are incomparable and differently sensitive, less experienced audiences cannot easily compare and interpret results, which is also prevalent in recent LLM benchmarks like Kola(Yu et al., 2023). Therefore, we utilized standardized global scores to evaluate the performance of LLMs on each task.
@@
where $\mu$(·) and $\sigma$(·) denote the mean and standard deviation.Next, we use the Min-Max scaling(Patro and Sahu, 2015) method to adjust the scores to the range of 0-100, making it easier to observe and compare the results. The final scores are presented as:
@@ 4.1 Experimental Setup
Using the GraCoRe benchmark, we investigate whether language models can understand and reason about graph structures through textual descriptions. Furthermore, we examine whether tailored prompts can improve their performance on graph-related tasks.
Models and Settings We evaluated a total of eight popular models on the GraCoRe benchmark, including four closed-source models and eight open-source models. The closed-source models are: OpenAI o1 This model is an LLMs with powerful reasoning capabilities launched by OpenAI on September 12, 2024. Due to time reasons, we only evaluated it in the main experiment., GPT-4o, GPT-4, and GPT-3.5(Achiam et al., 2023). The open-source models are: LLama3.1-ins-8b, LLama3-ins-8b, LLama2-7b-chat(Touvron et al., 2023), Chatglm3-6b, Chatglm2-32k-7b(Du et al., 2021), Vicuna-v1.5-7b, Vicuna-v1.5-16k-7b(Chiang et al., 2023) and Qwen2-7b-ins(Bai et al., 2023). More details about these models can be found in the Appendix B.
Figure 4: Effect of Graph Size.
@@ 4.2 Main Results
Task Dimensional Analysis Tables 3 and 4 present the performance of various LLMs across 19 tasks in the GraCoRe benchmark. Models generally performed better on graph understanding tasks, while graph reasoning proved more challenging. Closed-source models, particularly OpenAI o1, excelled in both areas, achieving the highest score of 1403.8, significantly surpassing other closed-source models. Among open-source models, Llama3.1-8b and Qwen2-7b-ins also performed well, ranking 5th and 7th with scores of 821.6 and 767.6, respectively. In contrast, Chatglm2-7b underperformed. The overall weaker performance on reasoning tasks was anticipated.
The average $z$-scores reveal a significant gap between open-source and commercial models. Only the Llama3.1-ins-8b and Llama3-ins-8b models achieve $z$-scores above 0, indicating above-average performance. This highlights the need for increased collaboration within the open-source community to enhance large language models. Notably, these two models demonstrate graph understanding and reasoning capabilities approaching those of GPT-3.5.
Ability Dimensional Analysis We further analyze Tables 3 and 4 to assess the models’ overall performance from a capability perspective, using radar charts to visually represent the second and third layers of the three-tier taxonomy. The left part of Figure 3 illustrates the performance of LLMs across four capability dimensions at the second layer, with each dimension’s score averaged across relevant tasks. Most models excel in graph understanding and semantic reasoning but require improvement in structural reasoning. The OpenAI o1 model notably addresses the structural reasoning weaknesses of other OpenAI models, offering a more balanced performance. However, most open-source models underperform in heterogeneous graph understanding tasks.
The right part of Figure 3 provides a detailed analysis of LLM performance across ten capability dimensions at the third layer. It reveals that large language models generally struggle with reasoning in graph theory, especially in complex tasks. However, models like OpenAI o1, GPT-4, and GPT-4o show strong and balanced graph processing abilities, excelling in more complex tasks. Notably, the OpenAI o1 model significantly outperforms others in graph theory reasoning, with performance several times better than other OpenAI models. In contrast, other models display inconsistent performance. For instance, Vicuna-v1.5-16k-7b excels in understanding heterogeneous graph structures but falls behind in other areas, suggesting that rich semantic information may enhance graph processing capabilities in certain contexts.
@@
Finally, we give the measured values of each model in each task in the main experiment in the Appendix D.
@@ 4.3 Further Analysis
Effect of Graph Size
Figure 4 shows the performance of four OpenAI LLMs on two graph comprehension and two graph reasoning tasks as node count increases. Results reveal a consistent performance decline across all tasks, particularly in the average degree task for graph comprehension, indicating increased computational complexity as node count grows. Despite this, the OpenAI o1 model shows strong computational ability, with minimal impact on comprehension tasks. However, both the shortest path and Hamiltonian path tasks in graph reasoning are notably affected, even for the robust OpenAI o1 and GPT-3.5 models. These findings indicate that current LLMs require further improvement in graph reasoning, particularly with large-scale graphs, where their performance drops considerably.
Effect of Random Sort Since our data consists of randomly generated graphs with nodes named by numbers, the performance of node-related tasks may be affected by changes in node order. In this study, we examine the impact of text order on the understanding of graph structures by large language models by comparing random sorting with sequential sorting. The results in Table 5 indicate that the model’s performance under sequential sorting is often superior to that under random sorting, particularly in graph path reasoning tasks, where the impact is significant. This suggests that renaming nodes and ordering them sequentially can enhance model performance in path reasoning problems. However, it also highlights the model’s lack of training on graph data with random sorting. Due to the inference cost, we chose only GPT-3.5 from OpenAI’s closed-source models for testing.
Model NN ST BR SP HP GPT-3.5 sort 0.993 0.954 0.504 0.117 0.243 random 0.968 0.865 0.346 0.070 0.048 Qwen2-7b sort 0.876 0.874 0.396 0.165 0.115 random 0.773 0.885 0.374 0.159 0.000 Llama3-ins-8b sort 0.889 0.735 0.200 0.189 0.285 random 0.786 0.809 0.200 0.198 0.059 Llama3.1-ins-8b sort 0.706 0.57 0.217 0.143 0.263 random sort 0.52 0.552 0.198 0.102 0.243 Chatglm3-6b sort 0.215 0.598 0.200 0.011 0.000 random sort 0.196 0.603 0.200 0.011 0.000 Chatglm2-32k-7b sort 0.008 0.005 0.63 0.039 0.000 random sort 0.011 0.004 0.504 0.043 0.000 Llama2-7b-chat-hf sort 0.075 0.616 0.200 0.159 0.002 random sort 0.105 0.413 0.200 0.133 0.002
@@
Figure 5: Effect of Text Enhancement.
Heterogeneous graphs provide rich semantic information that aids in understanding and reasoning about graph-related text. To assess whether LLMs can reason purely from structural data, we excluded enhanced text features, such as titles and abstracts, from the ACMText dataset. Figure 5 illustrates the performance of various models in a node classification task. Results show that GPT-4 and GPT-4o maintain strong prediction capabilities without text, indicating effective reasoning based solely on structure. However, GPT-3.5’s performance declines significantly without text, while open-source models show minimal impact from its absence.
@@ 5 Conclusion
This paper introduces GraCoRe, a benchmark designed to evaluate large language models’ (LLMs) ability to understand and reason with graph-structured data. We present a detailed, multi-level classification system for assessing model performance on graph-based tasks. Using GraCoRe, we evaluate 12 prominent LLMs, identifying significant limitations in their graph reasoning capabilities. Our experimental results validate the benchmark’s effectiveness in measuring LLM performance on graph tasks. Future research will focus on solving complex graph theory problems in large-scale pure graphs, with potential improvements in graph reasoning through methods like agents, chain-of-thought (CoT), and retrieval-augmented generation (RAG).
Limitations
@@
Acknowledgments
The research in this article is supported by the National Key R&D Pdrogram of China (Grant No. 2021ZD0112905), the National Science Foundation of China (Grant No. U22B2059, 62276083), the Major Key Project of PCL (Grant No. PCL2023A09), the Human-Machine Integrated Consultation System for Cardiovascular Diseases (Grant No. 2023A003). We also appreciate the support from China Mobile Group Heilongjiang Co., Ltd. @ on our research,the research is jointly completed by both parties.
Appendix A Comparison with existing benchmarks
In the Introduction, we outlined the advantages of the GraCoRe benchmark. First, it evaluates LLMs’ understanding and reasoning abilities on rich graph-structured data while also quantifying their performance on graph-related tasks. Second, GraCoRe tests a wider range of newer models and covers more tasks compared to existing benchmarks. Table 6 highlights the key differences between the GraCoRe benchmark and other benchmarks, demonstrating its broader scope and capabilities.
Appendix B Details on the data generation and models
@@
Our dataset consists primarily of two types of graph data: heterogeneous graphs and pure graphs. Based on these, we developed 19 tasks across 11 datasets to evaluate large models. This diversity ensures the comprehensiveness and robustness of our dataset.
@@
Figures 6 to 26 provide examples of prompts and graph data descriptions for each task. To ensure that generated results meet our capability and task-specific requirements, we concatenate the prompts for each task with the initial prompts.
We evaluated 12 of the latest LLMs, including OpenAI o1 reasoning model, launched on September 12, 2024. Table 7 presents more details on the models and their versions.
Appendix C More details about each task
Due to space limitations in the main text, this section will provide a detailed explanation of the specific details for each task. First, we will discuss the classification of model diagram understanding and reasoning capabilities associated with each task, and evaluate the number of test images used for each task. The evaluation results for all tasks are measured using accuracy (ACC). Table 10 summarizes the details of each task.
Appendix D Raw measurements of each task
In order to solve the problem of being unable to quantify the capabilities of each model in graph understanding and reasoning tasks, and to compare the differences in capabilities of each model, we use the standardized global scores metric to solve the above problem. However, we will still provide the actual measurement values of the large model in each task, and table 8 and table 9 shows the actual measurement values.
Benchmark Graph Type Evaluation Perspective Task # Model #
Pure Heterogeneous Task Model
NLGraph(Wang et al., 2024) 8 2
GPT4Graph(Guo et al., 2023) 10 4
GraphArena(Tang et al., 2024b) 10 10
GraphInstruct(Luo et al., 2024) 21 6
GraCoRe 19 12
Table 6: Differences between GraCoRe and existing benchmarks.
Table 7: More details about models.
Model Pure Graph Heterogeneous Graph
Attribute
Understanding
Graph
Memory
Graph
Recognition
Grap
Traversaling
Graph QA
and Quering
Subgraph
Extraction
NN(ACC) AD(ACC) CT(ACC) MS(ACC) TR(ACC) BR(ACC) BFS(ACC) NQ(ACC) RQ(ACC) RN(ACC) SE(ACC)
o1-mini 0.968 0.988 0.978 1.000 0.998 0.600 0.991 0.978 0.619 0.986 0.971
GPT-4o 0.993 0.706 0.753 0.999 0.630 0.339 0.809 0.745 0.562 0.955 0.713
GPT-4 0.991 0.626 0.946 0.998 0.578 0.211 0.809 0.843 0.560 1.000 0.725
GPT-3.5 0.993 0.490 0.698 0.954 0.459 0.504 0.898 0.489 0.291 0.637 0.439
Llama3.1-ins-8b 0.706 0.689 0.678 0.570 0.498 0.217 0.593 0.670 0.534 0.389 0.456
Llama3-ins-8b 0.889 0.289 0.555 0.735 0.563 0.200 0.896 0.295 0.556 0.892 0.275
Llama2-7b-chat 0.075 0.411 0.907 0.616 0.500 0.200 0.511 0.285 0.035 0.418 0.195
Chatglm3-6b 0.215 0.377 0.934 0.598 0.493 0.200 0.324 0.039 0.238 0.000 0.001
Chatglm2-32k-7b 0.008 0.255 0.446 0.005 0.515 0.630 0.057 0.568 0.238 0.000 0.125
Vicuna-v1.5-16k 0.397 0.328 0.210 0.601 0.472 0.335 0.248 0.377 0.542 0.802 0.204
Qwen2-7b-ins 0.876 0.165 0.958 0.874 0.520 0.396 0.343 0.454 0.556 0.079 0.635
Vicuna-v1.5-7b 0.110 0.415 0.934 0.007 0.459 0.200 0.139 0.016 0.558 0.281 0.032
Table 8: Performance of graph understanding.
Model Graph Structure Reasoning Graph Semantic Reasoning
Simple Graph
Theory Problems
Complex Graph
Theory Problems
Node Entity
Reasoning
Link Relationship
Reasoning
SP(ACC) MF(ACC) EP(ACC) HC(ACC) TSP(ACC) GC(ACC) NC(ACC) LP(ACC)
o1-mini 0.839 0.183 0.915 0.541 0.570 0.754 0.927 0.621
GPT-4o 0.346 0.020 0.783 0.15 0.043 0.363 0.929 0.737
GPT-4 0.228 0.015 0.791 0.167 0.035 0.320 0.912 0.674
GPT-3.5 0.117 0.017 0.772 0.243 0.028 0.448 0.796 0.568
Llama3.1-ins-8b 0.143 0.015 0.424 0.263 0.009 0.796 0.521 0.593
Llama3-ins-8b 0.189 0.022 0.359 0.285 0.015 0.489 0.459 0.649
Llama2-7b-chat 0.159 0.033 0.207 0.002 0.017 0.398 0.335 0.387
Chatglm3-6b 0.011 0.033 0.526 0.000 0.033 0.393 0.403 0.242
Chatglm2-32k-7b 0.039 0.033 0.324 0.000 0.028 0.198 0.212 0.222
Vicuna-v1.5-16k 0.133 0.013 0.215 0.000 0.015 0.128 0.100 0.377
Qwen2-7b-ins 0.165 0.026 0.480 0.115 0.048 0.265 0.468 0.617
Vicuna-v1.5-7b 0.180 0.015 0.222 0.002 0.026 0.233 0.383 0.014
Table 9: Performance of graph reasoning.
Level 3 Task Dataset Graph #
Attribute Understanding NN BG/TSG/GTG/SPG 1840
AD BG/TSG/GTG/SPG 1840
CT BG/TSG/GTG/SPG 1840
Graph Memory MS BG/TSG/GTG/SPG 1840
Graph Recognition TR Tree Structure Graph 460
BR Bipartite Graph 460
Graph Traversaling BFS Graph Traversal Graph 460
Graph QA and Querying NQ IMDBText 500
RQ IMDBText 500
RN IMDBText 500
Subgraph Extraction SE IMDBText 500
Simple Graph Theory Problems SP Shortest Path Graph 460
MF Max Flow Graph 460
EP Eulerian Graph 460
Complex Graph Theory Problems HC Hamiltonian Graph 460
TSP TSP Graph 460
GC Graph Coloring 460
Node Entity Reasoning NC ACMText 500
Link Relationship Reasoning LP ACMText 500
Table 10: Overview of Tasks and Datasets for Graph-based Problems.